# Information scrambling in quantum-walks: discrete-time formulation of Krylov complexity

### v2

@@ V Conclusion
In summary, the discrete-time quantum walks are quantum model which can be used to simulate a large number of phenomenon from quantum many-body physics as well as for construction of quantum algorithms. There experimental implementation on wide number of platform makes them suitable to study theoretical ideas. In this article, we studied information scrambling in discrete-time quantum walk using the out-of-time correlators and K-complexity as probe. The OTOC features the “shell-like” behavior, in which, at long-time limit, it goes to zero indicating no scrambling of the operator. The introduction of disorder (spatial or temporal) results in slowdown of information scrambling where the shape of light-cone confines up to localization length. The K-complexity shows a linear growth which ties to the fact that the operator at any time is approximately orthonormal to all operators at previous times. The disorder suppresses the K-complexity growth resulting its behavior of sub-linear. While the spatial disorder results in strong localization, the temporal disorder results in weak localization which is apparent from the behavior of both OTOC as well as K-complexity.
We would like to understand the effect of boundary conditions. In this work, we focused discrete-time quantum walk in infinite-dimensional 1D lattice, but from experimental point of view, it may be interesting to see the late-time behavior of these quantities $t>L$. In ref. [75], the information scrambling was studied on Clifford quantum cellular automata (QCA). These systems shown to break ergodicity, i.e., they exhibit quantum scarring. It was further shown that such a system could exhibit classical dynamics in some semi-classical limit. While the discrete-time quantum walks are not the same as QCA, but they can be regarded as the dynamics of the one-particle sector of a QCA [76]. Therefore, it be interesting if the connection between the two results can be made more direct. In this direction, one should note that Clifford QCAs model studied in ref. [75] also exhibits linear growth in K-complexity. It follows from the evolution of operator under Clifford QCAs in which a string of pauli operators maps to another. Since the pauli operator form a orthonormal basis, each evolution step generate a new basis element resulting in linear growth. While for infinite-dimensional case, this growth will persist forever due infinite dimensional Hilbert space, in case of finite-dimensional lattice, the growth will stop as the operator get backs to its initial state. Therefore, the linear growth in K-complexity follows by a sudden decay to zero followed by repetitive behavior which is refinance of recurrence dynamics.
@@
From the point of view of discrete-time quantum walks, which are the basis of many quantum algorithms, the study of information scrambling could provide the basis for designing algorithms that can efficiently explore the Hilbert spaces. In this context, the K-complexity which describe delocalization of operator in Hilbert space may be useful. Overall, the study of scrambling in both discrete-time quantum walks as well as in more generic systems with discrete-time evolution paves toward new physics in quantum many-body physics.
Acknowledgement: We wish to thank Aranya Bhattacharya for various useful discussions and comments about this and related works.

### v3

@@ V Conclusion
In summary, the discrete-time quantum walks are quantum model which can be used to simulate a large number of phenomenon from quantum many-body physics as well as for construction of quantum algorithms. There experimental implementation on wide number of platform makes them suitable to study theoretical ideas. In this article, we study information scrambling in discrete-time quantum walk using the out-of-time correlators and K-complexity as probe. The OTOC features the “shell-like” behavior, in which, at long-time limit, it goes to zero indicating no scrambling of the operator. The introduction of disorder (spatial or temporal) results in slowdown of information scrambling where the shape of light-cone confines up to localization length. The K-complexity shows a linear growth which ties to the fact that the operator at any time is approximately orthonormal to all operators at previous times. The disorder suppresses the K-complexity growth resulting in its sub-linear behavior. While the spatial disorder results in strong localization, the temporal disorder results in weak localization which is apparent from the behavior of both OTOC as well as K-complexity.
We would like to understand the effect of boundary conditions. In this work, we focused discrete-time quantum walk in infinite-dimensional 1D lattice, but from experimental point of view, it may be interesting to see the late-time behavior of these quantities $t>L$. In Ref. [75], the information scrambling was studied on Clifford quantum cellular automata (QCA). These systems shown to break ergodicity, i.e., they exhibit quantum scarring. It was further shown that such a system could exhibit classical dynamics in some semi-classical limit. While the discrete-time quantum walks are not the same as QCA, but they can be regarded as the dynamics of the one-particle sector of a QCA [76]. Therefore, it will be interesting to see if the connection between the two results can be made more direct. In this direction, one should note that Clifford QCAs model studied in ref. [75] also exhibits linear growth in K-complexity. It follows from the evolution of operator under Clifford QCAs in which a string of pauli operators maps to another. Since the pauli operator form a orthonormal basis, each evolution step generate a new basis element resulting in linear growth. While for infinite-dimensional case, this growth will persist forever due to the infinite dimensional Hilbert space, in case of finite-dimensional lattice, the growth will stop as the operator get backs to its initial state. Therefore, the linear growth in K-complexity follows by a sudden decay to zero followed by repetitive behavior which is the refinancing of recurrence dynamics.
@@
From the point of view of discrete-time quantum walks, which are the basis of many quantum algorithms, the study of information scrambling could provide the basis for designing algorithms that can efficiently explore the Hilbert spaces. In this context, the K-complexity which describe the delocalization of the operator in Hilbert space may be useful. Overall, the study of scrambling in both discrete-time quantum walks as well as in more generic systems with discrete-time evolution paves the way toward new physics in quantum many-body physics.
Acknowledgement: We wish to thank Aranya Bhattacharya for various useful discussions and comments about this and related works. We thank the anonymous referees for their thoughtful comments and constructive criticism.